### v1

# TerraMind: Large-Scale Generative Multimodality for Earth Observation

Abstract
We present TerraMind, the first any-to-any generative, multimodal foundation model for Earth observation (EO). Unlike other multimodal models, TerraMind is pretrained on dual-scale representations combining both token-level and pixel-level data across modalities. On a token level, TerraMind encodes high-level contextual information to learn cross-modal relationships, while on a pixel level, TerraMind leverages fine-grained representations to capture critical spatial nuances. We pretrained TerraMind on nine geospatial modalities of a global, large-scale dataset. In this paper, we demonstrate that (i) TerraMind’s dual-scale early fusion approach unlocks a range of zero-shot and few-shot applications for Earth observation, (ii) TerraMind introduces “thinking in modalities” (TiM)—the capability of generating additional artificial data during finetuning and inference to improve the model output—and (iii) TerraMind achieves beyond state-of-the-art performance in community-standard benchmarks for EO like PANGAEA. The pretraining dataset, the model weights, and our code will be open-sourced under a permissive license.
footnotetext: $\ast$ Equal contributionfootnotetext: $\dagger$ Equal supervision
1 Introduction
Earth observation (EO) increasingly benefits from multimodality because of the important integration of complementary information from different data sources. This becomes particularly relevant as data from specific satellite missions can be unavailable for a specific time or location due to low revisiting times or weather phenomena like cloud coverage. Vice versa, for the computer vision domain, EO data is an important playground for the development of new approaches as there is significant publicly available data of very high quality and complexity. The available modalities range from sensors of different satellite missions to relevant complementary information like digital elevation.
Figure 1: TerraMind represents the first any-to-any generative, and large-scale multimodal model for Earth observation pre-trained on 1 trillion tokens from global geospatial data. The model digests dual-scale representations at pixel-level and token-level simultaneously. TerraMindv1 outperforms other deep learning models for Earth observation in downstream applications and unlocks (i) generation, (ii) zero-shot and finetuning applications, and (iii) ”Thinking-in-Modalities” finetuning and inference.
In this work, we introduce TerraMind, the first large-scale, any-to-any generative, multimodal model for EO. With TerraMind, we introduce a dual-scale pretraining on pixel-level and token-level for generative multi-modality extending previous approaches like 4M [52] and demonstrate benefits over training primarily on tokens. TerraMind encodes high-level contextual information in tokens to enable correlation learning and scaling, while, additionally capturing important fine-grained representations using pixel-level inputs. During the pretraining, TerraMind predicts masked target tokens so that our pretraining objective boils down to a cross-modal patch classification problem that results in high-quality latent spaces. TerraMind is pretrained on a custom global-scale geospatial dataset named TerraMesh with 9 million samples that have been aligned spatiotemporally and across modalities [7]. In addition to radar and optical satellite imagery from the Copernicus Sentinel-1 (S-1) and Sentinel-2 (S-2) missions, our dataset contains task-specific modalities such as land use/land cover (LULC) and normalized difference vegetation index (NDVI) maps, metadata like digital elevation models (DEM) and geographic coordinates, and natural language in the form of captions. To the best of our knowledge, TerraMind represents the first truly generative, multimodal deep learning model for EO. We provide an overview on TerraMind’s performance in the community-standard benchmark PANGAEA [49] in Figure 2 and highlight the generative capabilities of TerraMind in Figure 3. In summary, our key contributions are as follows: (i) We introduce a dual-scale approach for generative multimodal pre-training leveraging data on pixel-level and token-level, which outperforms other fusion approaches, enables generation of artificial data, and facilitates zero-shot and few-shot applications. (ii) We introduce thinking in modalities – similar to chain-of-thought approaches in LLMs – for multi-modal models in EO, demonstrating that infusing generated data during finetuning improves the performance. (iii) We demonstrate that TerraMind outperforms other geospatial foundation models both in unimodal and multimodal settings.
Figure 2: TerraMind outperforms other geospatial foundation models on PANGAEA benchmark [49] in finetuning. Performance is measured in mIoU and min-max scaled per dataset with max performance displayed in parentheses.
Figure 3: Chained generation example of TerraMindv1-B starting from either optical, radar, or digital elevation data. Left is input, middle is artificially generated data by TerraMind, right represents the sensor measurements and ground truths, respectively. Note that TerraMind performance is upper-bounded by the tokenizer performance, especially with respect to the diffusion decoder.
2 Related Work
Computer vision in Earth observation. Computer vision (CV) has significantly advanced Earth observation (EO) [75]. Many CV techniques, originally developed for natural image processing, have been adapted to EO [62], often with minimal modifications. A wide range of tasks benefit from these methods, including classification [16], semantic segmentation [71] (e.g., land cover mapping [20, 21]), change detection [59] (e.g., disaster response [19]), object detection [39] (e.g., vessel identification [55]), and regression (e.g., biomass estimation [53]). Deep learning architectures like CNNs [74] and Vision Transformers (ViTs) [17] have demonstrated strong performance, often surpassing traditional remote sensing (RS) methods. However, EO presents unique challenges, including diverse sensor modalities [4] and geospatial heterogeneity [46]. An emerging paradigm in EO is self-supervised learning (SSL) [64] and geospatial foundation models (GFMs) [45], which aim to leverage vast amounts of unlabeled RS data to develop general purpose task models [32]. While off-the-shelf CV models have shown promising results [36], they do not fully exploit the unique characteristics of geospatial data. Many GFMs still rely on generic CV architectures [50], which were not explicitly designed to handle the complexities of EO, such as heterogeneous sensor sources (e.g., optical, radar, DEM) [29], integrated with auxiliary data (e.g., text) [42, 47], and expert knowledge (e.g., prioritizing specific bands or indexes). In this direction, TerraMind better integrates domain-specific properties, developing a customized and expandable multimodal learning strategy.
Multimodality in CV. The current paradigm of multimodal CV is driven by the integration of diverse data streams [69], such as natural images [73], natural language text [10], temporal video data [58], and weather [70], within large foundation models [8]. Starting from the alignment of images and texts [57], these models moved beyond simple feature extraction, towards nuanced contextual understanding. The ability to combine several modalities allows for unprecedented capabilities in complex tasks [30], evidenced by the rapid advancement of multimodal Large Language Models (MLLMs) [30], that excel in tasks such as scene understanding [12], visual question answering [18], and video analysis [24]. Recent advances in architectural design [31] and large scale pre-training [11] have enabled the development of models that learn highly effective cross-modal representations [41], which can then be adapted to a wide variety of downstream tasks [66].
Multimodality in EO. Multimodality in EO originates from data fusion and is typically understood as the integration of SAR and optical data [13, 28, 38, 25] or the combination of optical data with vector data [5]. Some studies have explored alternative combinations of data. In [15], the authors introduce a contrastive framework for comparing RS images and street views. Even different optical sensors can be considered different modalities [61, 48]. Similarly, several multi-view images (i.e. multimodal) datasets [44, 26, 54] are introduced. More recent approaches combined text and images [40], both for discriminative [42] and generative [34] purposes. Lately, different GFMs are trained in a multimodal way [4, 54, 68], still focusing either on a specific set of modalities (e.g., vision [68]) or tasks (e.g., generative [34]). To the best of our knowledge, no existing model has combined a wide and diverse amount of modalities both for discriminative and generative purposes, as TerraMind does. In Table 1, we provide a comparison of TerraMind with related model architectures.
Model Modalities Any-to-Any Generation Multi-Scale Features
RemoteCLIP optical, text
CROMA optical, radar
AnySat aerial, optical, radar, NAIP
DeCUR optical, radar
DiffusionSat optical, text
DOFA optical, radar, hyperspectral, NAIP
Galileo optical, radar, elevation, weather, location, population, …
TerraMind optical, radar, land use, elevation, vegetation index, location, text
Table 1: Comparison of TerraMind to other model architectures. While several model architectures have been developed to process multiple EO modalities, TerraMind represents the first truly generative model that can produce any modality from any other modality.
3 Dataset
For the pretraining of TerraMind and its tokenizers, we create a multimodal dataset called TerraMesh, which will be open-sourced to the community. TerraMesh builds on existing datasets, which we expand by adding modalities from external data sources or by applying pseudo-labeling. We provide an overview of the aligned image modalities and a detailed dataset description in the supplementary material.
Base datasets. TerraMesh is based on SSL4EO-S12 [65, 6] and MajorTOM-Core [23], two unlabeled remote sensing datasets containing co-aligned radar and optical imagery from Sentinel-1 and Sentinel-2 satellites. SSL4EO-S12 has lower geographic coverage but is multi-seasonal. MajorTOM-Core covers most of the Earth’s land surface at a single timestamp. For MajorTOM-Core, we apply a subsampling scheme based on LULC classes and ecoregions. TerraMesh includes a total of approximately 9 million globally distributed samples from both Sentinel-1 and Sentinel-2, each measuring 264$\times$264 pixels at 10m resolution.
Additional modalities. We obtain co-aligned yearly LULC maps by ESRIhttps://planetarycomputer.microsoft.com/dataset/io-lulc-annual-v02 with nine land use classes. Additionally, we leverage SEnSeI v2 [22] as a cloud and ice annotation model and update the ESRI LULC classes for better spatiotemporal alignment. NDVI maps are computed using the corresponding spectral bands from Sentinel-2. DEM is extracted from the Copernicus DEM 30m dataset [2], which provides global coverage of the Earth’s elevation at a 30m resolution. Captions are generated synthetically by constructing RGB images from Sentinel-2 patches using the corresponding spectral bands and processing them with LLaVA-Next [37]. A tailored prompt guides the model to describe the content of each image (see supplementary material for details). For geolocations, we round the latitude and longitude values from the center of each patch to the nearest quarter degree and save the resulting discretized coordinates as strings in a well-defined format (e.g., “lat=35.75 lon=-89.00”).
4 Methods
TerraMind pretraining is two-staged. We first pretrain unimodal tokenizer models, tokenize the modalities, and then leverage token-level and pixel-level input to pretrain the TerraMind encoder-decoder architecture. In the following, we describe those individual stages.
4.1 Tokenization
We develop modality-specific tokenizers to encode each modality into a sequence of discrete tokens for pretraining and decode token sequences back to their original forms. Thus, TerraMind is in principle compatible with any modality, as long as it can be tokenized and aligned with other modalities. For reasons of space, we delegate most experiments related to the tokenizer performances to the supplementary material.
Image-like modalities. We train autoencoder-based architectures with a quantization step in the bottleneck for image-like modalities such as Sentinel-1, Sentinel-2, LULC, NDVI, and DEM. Tokenizer encoders process an input image and generate a latent representation for each 16$\times$16 patch, which is then discretized with finite-scalar-quantization (FSQ) [51] into one of $N$ codewords, where $N$ represents the codebook size. All tokenizers use a vocabulary size of 16K besides the simpler LULC modality for which we use 4K. These codewords are then used by the diffusion decoder to reconstruct the original image. By mapping each codeword to a unique integer in $\{0,1,\dots,N-1\}$, we obtain discrete tokens for each image patch. We pretrain the tokenizers in a self-supervised setting. FSQ as quantization method enhances training stability [51] compared to vector quantization [63]. FSQ improves training stability by using a fixed codebook instead of a learnable one, eliminating the need for codebook-related loss terms. Notably, FSQ is heavily influenced by ideas of neural compression [27]. For our 12 channel Sentinel2-L2A data, we achieve compression rates of over 3000x by quantizing the data. We summarize the architecture of our tokenizers in Figure 4.
Figure 4: Tokenizer for image-like modalities combining finite-scalar quantization [51] with diffusion decoding.
Sequence-like modalities. We treat both captions and geolocations as text and use a single text tokenizer to process both modalities. By discretizing the geographic coordinates and representing them as strings in a structured format, we introduce special coordinate tokens into the tokenizer’s vocabulary. This approach allows us to encode geolocations as a sequence of discrete tokens, beginning with a latitude token followed by a longitude token. For the textual data, we modify the existing WordPiece tokenizer from BERT [33] rather than fitting a new tokenizer specifically for our captions since satellite imagery and corresponding captions have far less variations than natural images.
4.2 Pre-training
Architecture. TerraMind uses a symmetric Transformer-based encoder-decoder architecture proposed in [52], which is designed to process sequences of multimodal tokens. In addition to discrete tokens, our models also accept cross-modal pixel-level inputs, specifically satellite imagery and digital elevation maps. In contrast to [52], we leverage pixel-level data across a range of input modalities to introduce dual-scale feature representations in the input space, which help to structure the embedding space. For pixel-level inputs, we apply learnable patch-wise linear projections to generate patch embeddings for each 16$\times$16 patch, similar to the approach used in ViT [17]. We illustrate the pretraining paradigm in Figure 5 with an example of four modalities. The model is agnostic to processing tokens or patches in the input space, while the target is generally token-level data. In total, we use six pixel-level modalities and eight token-level modalities. Some of those modalities overlap, i.e. they are present both as pixel-level and token-level information which allows the model to learn a cross-modal patch-wise classification in one of the $N$ codebook classes.
Figure 5: Illustration of the pre-training task. Given a multimodal sample, random subsets of patches, input tokens, and target tokens are sampled from a Dirichlet distribution. Based on the encoded input, the decoder predicts target tokens for the masked input.
Masking strategy. TerraMind applies a masked modeling approach in the token space following [52]. The model leverages a set of randomly selected target tokens that have to be reconstructed from a randomly selected set of input tokens and pixel-level data. During pre-training, we sample input and target data from a Dirichlet distribution:
$f(\mathbf{x}\mid\boldsymbol{\alpha})=\frac{1}{B(\boldsymbol{\alpha})}\prod_{i=1}^{M}p_{i}^{\alpha_{i}-1}$ (1)
where $\mathbf{p}=(p_{1},p_{2},\dots,p_{M})$ is a vector of probabilities, and $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2},\dots,\alpha_{M})$ is a vector of weighting parameters that we leverage during pretraining as a steering wheel to guide the model. $B(\boldsymbol{\alpha})$ refers to a multivariate Beta function.
We opt for masked token reconstruction to familiarize the model with the absence of entire modalities, which is crucial for a high usability of a multimodal model in Earth observation. During pre-training, the model learns an internal representation of unseen modalities which is expected to benefit a range of downstream applications. In addition, sampling input and target tokens improves the computational efficiency of the pre-training, as each token is a compressed representation of a patch with compression factors of between 250x and 3000x depending on the modality. Finally, without tokenized representations of the image-like modalities, it is challenging to learn the correlation to sequence-like modalities. The overall training objective of TerraMind boils down to a cross-modal patch-level classification problem that we represent via a cross entropy loss:
$\mathcal{L}_{\text{CE}}=-\sum_{i=1}^{N}y_{i}\log(p_{i}),$ (2)
where $y_{i}$ is the one-hot encoded true class of token $i$, $p_{i}$ is the predicted probability for token $i$, $N$ is the total number of possible tokens. Interestingly, we can infer an upper bound loss for a random model. When a model outputs arbitrary target tokens, the cross entropy loss will collapse to the natural logarithm of the vocabulary size:
$\mathcal{L}_{\text{CE, random}}=-\sum_{i=1}^{N}y_{i}\log\left(\frac{1}{N}\right)=\log(N),$ (3)
which we validated empirically as an upper bound for the TerraMind pretraining loss. For a vocabulary size of 16K, a random model would have a validation loss of 9.68.
Scaling. We trained three versions of TerraMind scaling across model size, compute, and data. In addition, we pre-train different versions of TerraMind with respect to the number of dual-scale features. TerraMindv1-B is pre-trained on 500B tokens for 6 days on 32 NVIDIA A100 GPUs. The model uses dual-scale features from both token-level and pixel-level. During initial experiments, we observed significant improvements from scaling model size when switching from a tiny backbone to a small backbone to a base backbone. Therefore, we pre-trained TerraMindv1-L on a large backbone with 500B tokens on 32 NVIDIA A100GPUs trained for 10 days. Finally, to better understand the effect of scaling across the dual-scale feature representation, we pre-train TerraMindv1-B-single as a single-scale model on primarily token-level data with optical S-2 L2A data as only pixel-level input (compared to pixel-level S-2 L1C, S-2 RGB, S-1 GRD, S-1 RTC, and DEM in TerraMindv1-B and -L). TerraMindv1-B-single is pretrained on 500B tokens from over one million samples for 6 days on 32 NVIDIA A100 GPUs. We summarize the scaling behavior in model size, compute, and data in Figure 9 of the supplementary material. We additionally provide final validation losses in Table 8 comparing v1-B and v1-L with the theoretical random loss.
4.3 Generation
After pretraining, TerraMind is able to generate tokens for any modality, conditioned on any subset of input modalities. The generative capabilities also unlock various zero-shot tasks, such as water body segmentation, vegetation assessments, land use segmentation, image captioning, geolocation prediction, and text-to-image generation. For the generation of image-like modalities, the decoder receives mask tokens for the modality to be generated and predicts the corresponding tokens based on the processed input from the encoder. For sequence-like modalities, the decoder is initialized with a start-of-sequence token and generates the output autoregressively. After generating all output tokens, the corresponding tokenizer allows to map from token-space to image or text space. TerraMind additionally supports chained generation which ensures consistency across generated modalities. The chained generation represents a conditional probability distribution where the prior probability distribution $\mathbf{P(M_{1})}$ is given by the input modality $\mathbf{M_{1}}$, and all subsequent modalities $\mathbf{M_{2}}$ to $\mathbf{M_{K}}$ are generated based on their conditioning on the input modality and potential additional ones:
$P(\mathbf{M_{1}},\dots,\mathbf{M_{K}})=P(\mathbf{M_{1}})\prod_{i=2}^{K}P(\mathbf{M_{i}}\mid\mathbf{M_{1}},\dots,\mathbf{M_{i-1}}).$ (4)
5 Experiments
In this section, we describe the performance gains resulting from TerraMind and experiment with the unlocked capabilities through the any-to-any generation of TerraMind. We especially investigate the quality of the artificially generated data as an input to subsequent finetuning tasks similar to chain–of-thought in LLMs.
5.1 Foundational experiments
Multimodality vs. unimodality. As a first motivational experiment, we outline the benefit of leveraging multimodal data in Earth observation at the example of water body mapping. For this, we leverage the ViT-B encoders from the unimodal tokenizer models for S-1, S-2, and LULC. For fine-tuning, we concatenate the embeddings of the encoders, and train a segmentation head with four ConvNeXt [43] blocks (late fusion). The results in Table 2 (left) suggest that regardless of which modalities we combine, the combination of two modalities always outperforms each unimodal model. Combining all three modalities achieves the best overall performance.
Token-level fusion vs. late fusion. In Table 2 (right), we investigate the effects of fusing the inputs on a token level through masked token reconstruction compared to the late fusion from the previous paragraph. We observe that token-level fusion outperforms late fusion in water body mapping and the performance gains are particularly high when LULC data is not available. This suggests that early fusion captures an internal representation of the multimodal state—especially pronounced for LULC—that benefits fine-tuning. With those findings in mind, we will explore the effects of using additional multi-modal pixel-level input in a dual-scale pretraining in Section 5.5.
Input Late fusion Token-level fusion
S-1 61.01 63.94 (2.93$\uparrow$)
S-2 72.70 76.32 (3.62$\uparrow$)
LULC 71.77 70.96 (0.81$\downarrow$)
S-1 + S-2 73.83 76.74 (2.91$\uparrow$)
S-1 + LULC 73.86 73.76 (0.10$\downarrow$)
S-2 + LULC 75.65 77.04 (1.39$\uparrow$)
S-1 + S-2 + LULC 76.00 76.88 (0.88$\uparrow$)
Table 2: Task: Water body mapping on Sen1Floods11 [9], performance measured in IoU on water class. Model sizes and architectures are comparable. Left column: Late fusion of tokenizers. Effect of leveraging multimodal data in late-fusion fine-tuning of frozen modality-specific ViT-B models. The average improvement of full multimodality over the individual unimodal performance is 7.5pp IoU. Right column: Finetuning results of TerraMind.v1-B-single as a mid fusion approach based on masked correlation learning. Performance gains over late fusion are depicted as percentage points increases/decreases in parentheses.
5.2 Generation experiments
TerraMind supports the generation of any modality from any other modality. In the following, we provide examples of the generation performance starting from: (i) an information-rich modality, like optical S-2 L2A data, and (ii) minimal information based on the geolocation. In Figure 3, we observe that TerraMind performs very well in generating image-like modalities like radar data, land use maps, and digital elevation maps from optical S-2 L2A data. Interestingly, TerraMind is also able to geolocate the optical image with a low deviation of less than 0.5 degree latitude/longitude. We provide additional experiments on the accuracy of geo-localization prediction in Section 5.3.
Figure 6: Randomly selected chained generation example with uni-modal geo-location input data. Top row is artificially generated data by TerraMind, buttom row represents a ground truth sample at this grid location, respectively.
While we expect to see accurate generations from information-rich modalities like optical data, it is particularly interesting to understand how TerraMind deals with low information content. Therefore, we prompt TerraMind to generate a subset of modalities starting from the geolocation in Figure 6. Interestingly, for a geolocation from the middle-east, the model generates an optical image that resembles a desert. While the generated optical image is based on the right context, the actual structure is unsurprisingly different from the ground truth. Based on the chained generation, this difference ripples down across all other modalities as well causing consistent but inaccurate generations. This example emphasizes the relevance of access to information-rich, fine-grained features to facilitate accurate generations.
5.3 Zero-shot experiments
Based on its generative capabilities, TerraMind unlocks several zero-shot applications, like land-use segmentation, water body mapping, geo-localization, and vegetation mapping. In the following, we focus on water body mapping and geo-localization as image- and sequence-level zero-shot tasks.
Water body mapping. In Table 3, we compare the zero-shot performance of TerraMind with its fine-tuned performance and other finetuned benchmarks for water body mapping. Overall, TerraMindv1-B-single achieves a zero-shot IoU of 45.4% compared to SOTA-level fine-tuning performance of 84.4% of DeCUR. In ablations with TerraMindv1-single trained on DynamicWorld LULC data, we boost this to up to 69.8% suggesting that TerraMind harnesses up to over 80% of the SOTA performance in zero-shot setting. Additionally, it’s notable that none of the benchmarking model can be applied in a zero-shot context, highlighting the relevance of TerraMind’s capabilities.
Model Input Type IoUWater
TerraMindv1-B S-2 zero-shot 45.40
TerraMindv1-B-single S-2 zero-shot 69.75
Prithvi 2.0 / DeCUR / … zero-shot N/A
Baseline [9] S-2 finetune 31.25
Prithvi 2.0 300M S-2 finetune 80.97
DeCUR S-2 finetune 82.17
Table 3: Zero-shot results of TerraMind on water body mapping compared to fine-tuned performance of benchmarks.
Geo-localization. TerraMind has strong capabilities in predicting the geolocation of a specific data instance. To better visualize the geo-localization capabilities, we prompt the model for the most likely locations of the land use class “bare land” (deserts etc.) in a Monte-Carlo-sampling in Figure 7(a). The probability distribution of the model fits the expectation of where to find bare land accurately, highlighting the Sahara region and middle-east, as well as Mexico and Southern California.
(a) Prediction distribution of the land use class “bare land” with a sampling temperature of $T=1.0$. TerraMind has an accurate internal representation of the geolocation of specific contexts, like land use classes.
(b) Distribution of predicted geo-locations for an optical S-2 L2A sample from the validation set. $\bullet$ is the correct location, $\bullet$ are Monte-Carlo sampled locations from TerraMind, $\bullet$ represents the distribution of training locations. TerraMind’s geo-localization seems to be based on similar optical samples in the training dataset for which TerraMind then outputs the geolocation.
Figure 7: Monte-Carlo sampling approach for TerraMind’s geo-location prediction.
To better understand how TerraMind assigns geolocations to samples, we further employ a Monte-Carlo sampling on the latitude-longitude grid for an optical tile from the validation data in Figure 7(b). We observe that while TerraMind is not predicting the correct geolocation ($\bullet$), there is a very high likelihood that the predicted geolocation is one of the adjacent grid points ($\bullet$). Thus, TerraMind likely looks out for similar looking optical samples in the training data ($\bullet$) and returns their geolocation.
5.4 Few-shot experiments
Model Input EuroSAT METER-ML
CLIP-ViT-B/16 S-2 RGB 52.39 28.13
CLIP-ViT-B/16 NAIP 31.73
DeCUR S-2 L1C 50.54 27.87
Prithvi 1.0 100M S-2 L1C 60.11 26.08
Prithvi 2.0 300M S-2 L1C 61.06 28.26
Terramindv1-B S-2 L1C 70.83 33.90
Terramindv1-B NAIP 32.23
Table 4: 1-shot 5-way classification results on EuroSAT and METER-ML measured in mean accuracy $\uparrow$, averaged over 200 runs. TerraMind outperforms benchmarks from CV and EO domain, suggesting a well-structured latent space.
TerraMind is pretrained on a cross-modal patch classification objective. Therefore, we expect a well-structured latent space that clusters different concepts accurately. To investigate our hypothesis, we apply 1-Nearest-Neighbor (1-NN) classification experiments in the community-standard setting of 1-shot 5-way on two datasets: EuroSAT and METER-ML. In those experiments, there are no weight updates of any kind, so that we can assess the quality of the embedding space structure.
In Table 4, we observe that TerraMind outperforms several other benchmarks from both the CV and EO domain on the EuroSAT dataset by at least 10pp in accuracy. Our results further show that for methane source classification on METER-ML, TerraMind outperforms benchmark models and generalizes to high-resolution NAIP data with one order of magnitude higher resolution than the pre-training data. We present additional experiments with other few-shot settings in the supplementary material.
5.5 Fine-tuning experiments
Besides the novel capabilities that TerraMind introduces, we benchmark the fine-tuning performance of TerraMind in both unimodal and multimodal settings following the community-standard PANGAEA benchmark [49]. We summarize the results in Table 5. Overall, TerraMindv1-B outperforms all other GeoFMs by at least 3pp avg. mIoU. Importantly, we report additional results in Table 11 of the supplementary material that show how TerraMind is the only foundation model approach in EO that outperforms task-specific U-Net models across the PANGAEA benchmark. Performance is approximately 2pp avg. mIoU higher for Terramindv1-L, with a peak of 5pp in multimodal datasets. Furthermore, TerraMindv1-L outperforms also specialised ViT baselines by 5pp avg. mIoU. Finally, TerraMindv1-B outperforms TerraMindv1-B-single by 1pp. avg. mIoU. The latter result demonstrates the benefit of using dual-scale features when comparing v1-B-single with v1-B. Note that per suggestion of the PANGAEA authors, we exclude the xView2 and BioMassters task as we could not reproduce the reported performances.
Model BurnSr MADOS PASTIS Sen1Fl11 FBP DEN CTM-SS SN7 AI4Farms Avg. mIoU $\uparrow$ Avg. Rank $\downarrow$ CROMA 82.42 67.55 32.32 90.89 51.83 38.29 49.38 59.28 25.65 55.29 5.06 DOFA 80.63 59.58 30.02 89.37 43.18 39.29 51.33 61.84 27.07 53.59 6.33 GFM-Swin 76.90 64.71 21.24 72.60 67.18 34.09 46.98 60.89 27.19 52.42 8.00 Prithvi 1.0 100M 83.62 49.98 33.93 90.37 46.81 27.86 43.07 56.54 26.86 51.00 8.89 RemoteCLIP 76.59 60.00 18.23 74.26 69.19 31.78 52.05 57.76 25.12 51.66 9.22 SatlasNet 79.96 55.86 17.51 90.30 50.97 36.31 46.97 61.88 25.13 51.65 8.56 Scale-MAE 76.68 57.32 24.55 74.13 67.19 35.11 25.42 62.96 21.47 49.43 9.44 SpectralGPT 80.47 57.99 35.44 89.07 33.42 37.85 46.95 58.86 26.75 51.87 8.11 S.-S12-MoCo 81.58 51.76 34.49 89.26 53.02 35.44 48.58 57.64 25.38 53.02 8.00 S.-S12-DINO 81.72 49.37 36.18 88.61 51.15 34.81 48.66 56.47 25.62 52.51 8.67 S.-S12-MAE 81.91 49.90 32.03 87.79 51.92 34.08 45.80 57.13 24.69 51.69 10.28 S.-S12-Data2Vec 81.91 44.36 34.32 88.15 48.82 35.90 54.03 58.23 24.23 52.22 8.72 TerraMindv1-B 82.42 69.52 40.51 90.62 59.72 37.87 55.80 60.61 28.12 58.35 2.94 TerraMindv1-L 82.93 75.57 43.13 90.78 63.38 37.89 55.04 59.98 27.47 59.57 2.56
Table 5: Performance evaluation of frozen TerraMind encoder with trainable UperNet head following the PANGAEA evaluation protocol across nine benchmark datasets. We report the mIoU (%) $\uparrow$ and highlight the best model in bold and the second best is underscored.
Finally, we assess the impact of leveraging multimodal data as input to TerraMindv1-B compared to utilizing either optical or radar data as unimodal input to better understand the effect of leveraging multimodal data in finetuning. We observe that across all three multimodal tasks, TerraMindv1-B performs best with access to both optical and radar data.
PASTIS Sen1Fl11 CTM-SS
S-1 20.04 80.39 24.45
S-2 40.20 89.57 50.90
Multimodal 40.51 90.62 55.80
Table 6: Benefit of using multimodal input in TerraMind.v1-B on multimodal datasets of PANGAEA. We report the mIoU (%) $\uparrow$.
5.6 Thinking in modalities
Based on the generative capabilities of TerraMind and the fact that multi-modal data improved downstream task performances in many scenarios, we additionally investigate the value of artificially generated data for finetuning tasks. Borrowing on the idea of chain-of-thought finetuning of large language models, we use TerraMind to generate artificial data as an intermediate step to improve finetuning. We provide an overview on this new method in the beginning of this paper in Figure 1.
Fine-Tuning Input IoUWater mIoU
TerraMindv1-B S-1 68.00 81.06
TerraMindv1-B S-2 82.26 89.70
TerraMindv1-B TiM S-1 + gen. LULC 72.25 83.65
TerraMindv1-B TiM S-2 + gen. LULC 84.75 91.14
Table 7: Thinking-in-modalities (TiM) tuning compared with standard full fine-tuning approaches on the Sen1Floods11 dataset.
We employ the TiM tuning approach on water body mapping and during finetuning use radar or optical data to generate additional artificial LULC data which we leverage along the input. Our results in Table 7 indicate a superior performance of the TiM finetuning compared to leveraging uni-modal data by up to 2pp mIoU. This finding points us in the direction of TerraMind being able to generate data that improve finetuning to downstream applications.
6 Conclusion
TerraMind’s pre-training approach of combining token-level and pixel-level information has unlocked a range of new model capabilities in Earth Observation. TerraMind demonstrates not only beyond state-of-the-art performance when tested against 12 other geospatial models on the PANGEA benchmark, it also represents the first truly generative multimodal model in the domain. Because of the ability of integrating heterogeneous and diverse data sources demonstrated by the Terramind’s approach, we expect that TerraMind-like models will expand soon to multi-temporal, multi-resolution, and hyperspectral data. This will allow to maximise impact in real world applications and fully leverage the data rich ecosystem available in the Earth Observation domain.
7 Acknowledgements
This work is part of the Fostering Advancements in Foundation Models via Unsupervised and Self-supervised Learning for Downstream Tasks in Earth Observation (FAST-EO) project, which is funded by the European Space Agency (ESA) $\Phi$-lab under contract No. 4000143501/23/I-DT. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS [3] at Jülich Supercomputing Centre (JSC).
Supplementary Material
In the following, we provide additional information on our data, the pretraining of TerraMind and its tokenizers, the quality of the tokenization, any-to-any generation matrices, and comparisons of TerraMind in unimodal and multimodal finetuning against specialized U-Net and ViT models.
8 TerraMesh Dataset
All versions of TerraMind have been pretrained on TerraMesh or a subset of it. TerraMesh is a comprehensive multimodal Earth observation dataset designed for large-scale model pre-training. It will be made publicly available under a permissive license in a preprint during the review process of this paper. The dataset includes nine modalities and we visualize examples of the dataset in Figure 8.
The dataset contains over 9 million globally distributed, spatiotemporally aligned samples across nine core modalities. Each modality is precisely co-registered at a 10-meter resolution, primarily based on Sentinel-2 grids. The S-1 and S-2 samples are sourced from MajorTOM-Core [23] and SSL4EO-S12 v1.1 [6]. It integrates Sentinel-1 SAR data with Sentinel-2 optical data (L1C top-of-atmosphere and L2A bottom-of-atmosphere reflectance), ensuring versatility for various downstream tasks. Because the source datasets contain only one S-1 product, each sample has either S-1 GRD or S-1 RTC data. Additionally, TerraMesh includes normalized difference vegetation index (NDVI) maps derived from Sentinel-2, Copernicus digital elevation model (DEM) data providing topographic context, and land-use/land-cover (LULC) maps from ESRI, enhanced with accurate cloud masks generated by the SEnSeI v2 model[22].
To ensure broad geographic and thematic diversity, TerraMesh employs subsampling techniques, selectively including representative samples from each global ecoregion and land-cover class, while downsampling highly homogeneous regions such as deserts and tundra. Another critical aspect is the data preprocessing pipeline, which includes reprojection, temporal alignment, and filtering to minimize missing data and artifacts, ensuring high-quality, analysis-ready samples.
Figure 8: Visualization of the spatial-temporal alignment across modalities in TerraMesh. S-2 L2A uses IRRG pseudo-coloring and S-1 RTC is visualized as VH-VV-VH. To account for their value ranges, VV is scaled from -30 – 5 dB while VH is scaled from -40 – 0 dB. Copernicus DEM is scaled based on the image value range.
TerraMind.v1-B-single was pre-trained on a subset of TerraMesh with one million samples, specifically the SSL4EOS12 v1.1 locations, using only four image modalities: S-2 L2A, S-1 GRD, DEM, and LULC. Additionally, we performed continuous pre-training with image captions. These captions were created using LLaVA-Next [37] and Overture Maps data. The automated captioning pipeline includes a prompt with a chain-of-thought process to generate diverse captions. The captioning model is asked to generate three question-answer pairs and describe the full image later. We use the S-2 RGB bands and Overture base layer tags as inputs. Domain experts evaluated a subset of 1.3k captions, resulting in 69% of the captions without any hallucinations while the average completeness scores were 3.87 on a scale from 0 to 5. The captions will be released as an additional preprint during the review process of this paper.
9 Pretraining details
In this section, we give additional details on the pretraining of both TerraMind and its tokenizers.
9.1 Tokenizer models
We pretrained the TerraMind tokenizers for image-like modalities with DDP on 4 GPUs for a total of 100 epochs on the respective modality of TerraMesh. We use a base learning rate of 1e-4, an effective batch size of 64 samples per GPU, i.e. the global batch size is 256. We reach a GPU utilization of 99% for single channel modalities like LULC and NDVI, and over 80% for all multi-channel modalities.
9.2 TerraMind
We pretrained both TerraMindv1-B and TerraMindv1-L with DDP on 32 GPUs. We determine the global batch size based on initial experimental runs comparing a global batch size of 2K, 4K, and 8K. In addition, we determine the base learning rate starting from 1e-4 and iteratively experimented with half and double learning rates. Ultimately, we end up with a base learning rate of 2e-4 for a cosine annealing scheduler set to run for 500B tokens. For the v1-L model, we reach a GPU utilization of 85+%. Overall, the training of TerraMindv1-B took 12 days on 32 A100 GPUs, i.e., 9’216 GPU hours. Over the course of the pretraining, we also experiment with different configurations of the Dirichlet sampling distribution. In total, the pretraining experiments have been approximately three times larger than the final runs resulting in approximately 30K GPU hours allocated for pretraining.
We provide an overview on the scaling dynamics when going from TerraMindv1-B to TerraMind v1-L in Figure 9 with identical hyperparameters and compute. Overall, as expected, we observe a significant gap in the validation losses across modalities. We finally provide the validation losses per modality after pretraining of TerraMindv1-B and TerraMindv1-L in Table 8.
Figure 9: Example of the scaling behavior of TerraMind comparing v1-B and v1-L models for the first 350B tokens on the validation loss of optical S-2 L2A data. Overall, TerraMind-L outperforms TerraMind-B after approximately 10% of the training schedule of the large model.
Model S-2 L2A S-1 GRD S-1 RTC DEM NDVI
Random 9.68 9.68 9.68 9.68 9.68
V1-B 5.67 7.84 7.64 2.19 6.42
V1-L 5.34 7.69 7.53 2.14 6.25
Table 8: Validation losses of full pre-training of TerraMindv1-B and v1-L.
10 Tokenizer performance and general learnings
In the following, we provide details on the tokenizations of TerraMind. At least for image-like modalities, the tokenizations represent an important and computationally heavy phase of the pretraining, which is why we highlight important learnings in the following.
Learnings. Overall, we learned that the tokenizer performance can be quite sensitive, which is especially related to the significant bottleneck compression of up to 3000x after the encoder. When leveraging finite-scalar quantization (FSQ) instead of vector quantization (VQ), we observed exactly what the original FSQ paper [51] claims: FSQ makes quantization easier – yet in our experiments it did not improve the reconstruction performance in terms of MSE losses. We leverage FSQ as the training was more stable and less sensitive to the learning rate, which is likely related to the fact that, unlike VQ, FSQ does not require an additional codebook loss. We still observed that all tokenizer models were sensitive to the learning rate, with higher learning rates resulting in non-differentiability (NaN losses), and low learning rates caused blurry results.
In addition, we experimented with the codebook size. In our experiments, we observed that the level of detail in the reconstructions was significantly higher for single channel input compared to multi channel input (e.g., 12 band S2-L2A data). Naturally, with less channels, the compression bottleneck for equal-sized codebooks is lower. Therefore, we hypothesized whether multi-spectral data requires larger codebook sizes to obtain higher level of detail in the reconstructions. In contrast to our expectation, when increasing the codebook size over 16K for modalities with more than three input channels, the reconstructions had significant artefacts. This suggests that even though the compression bottleneck is lower, higher codebook sizes are more difficult for the model to use, which is in line with previous literature. However, we were surprised to see more artefacts in the reconstructions of models with a codebook size 32K compared to 16K.
Finally, we experimented with exponential moving average (EMA) updates for the tokenizer models. As expected, the models were less responsive to gradient updates. The resulting reconstructions smoothed out more of finegrained features. Together with the generative diffusion process in the tokenizer decoder, the resulting reconstructions often looked like hallucinations, e.g. bridges over rivers were not existing anymore in the reconstruction images. We therefore decided to ommit expotential moving average in our tokenizer models.
Performance. In the following, we assess the accuracy of our tokenizer models. Besides visual quality assessments and quantitative assessments with MSE metrics, we were particularly interested in whether our tokenizers exhibit geospatial biases. Understanding this is crucial to ensure TerraMind has a uniform level of performance across the globe. In addition, we investigate the reconstructions of radar data in more detail, as radar data by nature includes significant noise in the amplitude data. This could interfere with the noise generation in the diffusion process of the decoder, which is why we assess the structure of the reconstructions using SSIM and PSNR metrics.
In Figures 10 to 12, we provide an overview on the spatial distributions of the S-1 GRD, S-2 L2A, and DEM tokenizer on the validation data of the SSL4EO-S12 subset which is focused on urban areas and therefore relevant for many downstream applications. Overall, we observe low MSE errors and particularly low deviation across geographic regions. For optical S-2 data, we observe minor difficulties in reconstructing images from Northern Asia, which we manually investigated. Overall, the vast majority of those samples are depicting snowy/icy conditions that have very high reflectance values of up to 12,000 compared to a normal range of [0, 255] in RGB data. On those long tail distribution samples, the S-2 tokenizer naturally has more difficulties.
Figure 10: Spatial distribution of mean squared errors of the S-1 tokenizer on the validation set of the pretraining data.
Figure 11: Spatial distribution of mean squared errors of the S-2 tokenizer on the validation set of the pretraining data.
Figure 12: Spatial distribution of mean squared errors of the DEM tokenizer on the validation set of the pretraining data.
S1-tokenizer quantitative analyses. In the following, we pay particular attention to the performance of the radar S-1 tokenizer, which might be more challenging to train on a reconstruction task due to the inherent speckle noise in radar satellite data. We therefore evaluate the reconstructions of the S-1 tokenizer using the structural similarity index (SSIM) and peak signal-to-noise ratio (PSNR). Both input and reconstruction for S-1 are in a dB scale. In addition to S-1 evaluation metrics being computed in the dB space in Table 9, they also are calculated in the denormalized space. On the contrary, the S-2 evaluation metrics are computed in the normalized space.
We give a more extensive background on radar data in the following for interested readers and non-EO experts. Reconstructing realistic and accurate synthetic aperture radar (SAR) S-1 VV and VH data is challenging due to factors inherent in the specific characteristics of SAR and the S-1 mission. SAR data is affected by complex interactions between the radar signal and Earth’s surface. SAR is based on radar backscatter, which is influenced by surface roughness and moisture content. The interaction of radar waves with different surfaces, including vegetation structure and urban environments, can produce complex backscatter patterns. The two polarizations, VV and VH, capture different scattering mechanisms: VV is sensitive to surface roughness and vegetation, while VH captures cross-polarized interactions that are influenced by surface and volumetric features [14, 56, 35]. In addition, SAR inherently contains speckle noise, which obscures fine details, making it difficult to extract accurate information. To evaluate the SAR data tokenizers of TerraMind, we employ various evaluation metrics to assess quality and accuracy. We compute the MSE for quantifying pixel-level differences, the SSIM to compare image structural content, and the PSNR [72, 67, 1].
We present the results in Table 9. Overall, we observe high to very high values in both SSIM (generally above 90% SSIM) and PSNR (generally above 55dB). These results validate the high quality of the encoding and decoding of the S-1 data using our tokenizers. Furthermore, we observe that the reconstructions are slightly better for the VV component of S-1 compared to the VH component. For reasons of comparison, we additionally provide the results for the S-2 tokenizer, which exhibits slightly lower SSIM score and strong PSNR scores. We additionally perform qualitative evaluations by visual inspection to identify artifacts, noise and other inconsistencies that may not be captured by numerical evaluation metrics on their own.
In addition to MSE, SSIM and PSNR in Table 9, for TerraMindv1-B, for the tokenizers GRD VV and VH, the correlation between the estimated and ground truth intensity values is $0.625$ and $0.701$, respectively, while for TerraMindv1-B-single, the correlations are $0.576$ and $0.670$. This improvement is likely related to the fact that the v1-B tokenizers have been trained on more (i.e. in number) and also much more diverse data. Moreover, the correlation between MSE and SSIM for all the test data samples is $-0.613$, which is interesting as SSIM is a different metric from MSE, while the correlation between MSE and PSNR is $-0.928$, i.e. closer to $-1$. These correlations are for S-1 GRD VV, while for VH, the correlations are $-0.416$ and $-0.929$.
TerraMindv1-B TerraMindv1-B-single
Tokenizer MSE SSIM PSNR MSE SSIM PSNR
S-1 GRD VV 0.118 0.961 58.1 0.154 0.961 57.8
S-1 GRD VH 0.172 0.940 56.2 0.207 0.939 56.0
S-2 L2A 0.037 0.926 66.3 0.032 0.933 66.8
Table 9: Evaluation of SAR VV and VH and S-2 reconstructions by the TerraMind tokenizers using MSE $\downarrow$, SSIM $\uparrow$ and PSNR $\uparrow$ on the validation dataset of the SSL4EO-S12 subset (8.5k samples).
11 Any-to-any generation
In Figure 13, we provide an example of any-to-any generation on four image-like modalities and two sequence-like modalities. Overall, we observe that when we start from modalities with high information content (e.g., fine-grained image-like modalities), the reconstructions are particularly good. Even with less information content, the model is able to generate consistent artificial data. However, we can clearly observe that the quality compared to the ground truth (represented by the input in the left of the figure) is decreasing. Finally, it is interesting to see how artefacts are introduced by the model when starting from lower information content in the input. For example, when promting TerraMind to generate data from DEM input, we observe that the model pays significant attention to the darker streams in the DEM image, which are later generated as a river in LULC.
Figure 13: Any-to-any generation example of TerraMindv1-B-single. Fine-grained input like optical and radar achieve particularly good performances.
12 Additional experiments
In the following, we provide additional experiments, especially with regard to the quality of the latent space and the full finetuning performance. To understand the quality of the latent space, we compute performances of nearest neighbor approaches for image classification tasks or using prototypical neural networks. We assess the performance of full finetuning by comparing with end-to-end trained, task-specific models like U-Nets and ViTs. We additionally compare the quality of the generations with the pseudo-labels used to pretrain TerraMind in an ablation experiment in a zero-shot setup.
12.1 Few-shot experiments
We present additional few-shot experiments with the EuroSAT and METER-ML dataset in Table 10. We use the embeddings of the pre-trained encoders without any additional fine-tuning. The patch embeddings of each image are averaged for image-level classification tasks.
The experiments include four different few-shot settings with varying numbers of examples and classes. 5-way refers to sampling five classes per run, while full-way describes experiments with all dataset classes per run. 1-shot and 5-shot indicate that one or five images are sampled for each class per run. 5-shot experiments with five support samples per class are using Prototypical Networks [60] for classification. This approach averages the embeddings of the selected labeled images (support set) and classifies the target images (query set) based on the class prototype with the lowest Euclidean distance from each sample. In the 1-shot setting, Prototypical Networks are mathematically equal to 1-Nearest-Neighbor classification. We refer to the original paper for details [60]. Different from literature, we evaluate each run on the full test set instead of subsampling query images.
TerraMind performs best on both datasets, outperforming all other geospatial foundation models as well as the CLIP vision encoder [57]. Interestingly, the base version leads to overall better results than the large model. Similarly, Prithvi’s smaller 1.0 version has comparable results to its larger 2.0 300M version, indicating that model size has only a limited effect on few-shot performance.
In addition to S-2 L1C, the METER-ML dataset provides high resolution RGB images from NAIP with 1 m resolution. Only CLIP and TerraMind can process RGB images without any fine-tuning. While CLIP profits largely from the higher resolution inputs, TerraMind only performs marginally better and sometimes worse than with multispectral S-2 data. Notice that TerraMind shows similar performance gaps as CLIP when comparing NAIP data to S-2 RGB. This indicates that additional multispectral channels have a comparable effect on few-shot performance as high-resolution images.
EuroSAT METER-ML
Model Input 5-way 1-shot 5-way 5-shot full-way 1-shot full-way 5-shot 5-way 1-shot 5-way 5-shot full-way 1-shot full-way 5-shot
CLIP-ViT-B/16 S-2 RGB 52.39 74.39 39.71 62.80 28.13 37.51 22.21 30.92
CLIP-ViT-B/16 NAIP 31.73 42.84 24.86 36.41
DeCUR S-2 L1C 50.54 64.35 37.53 50.82 27.87 33.64 20.95 27.21
Prithvi 1.0 100M S-2 L1C 60.11 73.29 46.86 60.66 26.08 35.81 22.33 29.21
Prithvi 2.0 300M S-2 L1C 61.06 73.21 47.47 60.47 28.26 36.13 22.52 29.59
TerraMindv1-B S-2 L1C 70.83 87.94 57.48 79.66 33.90 43.89 26.85 37.41
TerraMindv1-B NAIP 32.23 44.75 25.53 37.85
TerraMindv1-L S-2 L1C 70.07 86.29 56.58 77.39 33.09 42.72 26.02 36.34
TerraMindv1-L NAIP 32.59 44.99 25.94 38.29
Table 10: Few-shot classification results on EuroSAT and METER-ML measured in mean accuracy $\uparrow$ averaged over 200 runs. 5-way refers to five randomly sampled classes per run, which is a default setting used in few-shot learning. Full-way refers to sampling all dataset classes, i.e., ten EuroSAT classes and seven METER-ML classes. We highlight the best two models in bold and underlined.
12.2 Finetuning comparisons with baseline models
Model BurnSr MADOS PASTIS Sen1Fl11 FBP DEN CTM-SS SN7 AI4Farms Avg. mIoU Avg. Rank CROMA 82.42 67.55 32.32 90.89 51.83 38.29 49.38 59.28 25.65 55.29 6.61 DOFA 80.63 59.58 30.02 89.37 43.18 39.29 51.33 61.84 27.07 53.59 8.22 GFM-Swin 76.90 64.71 21.24 72.60 67.18 34.09 46.98 60.89 27.19 52.42 10.00 Prithvi 1.0 100M 83.62 49.98 33.93 90.37 46.81 27.86 43.07 56.54 26.86 51.00 11.00 RemoteCLIP 76.59 60.00 18.23 74.26 69.19 31.78 52.05 57.76 25.12 51.66 11.22 SatlasNet 79.96 55.86 17.51 90.30 50.97 36.31 46.97 61.88 25.13 51.65 10.67 Scale-MAE 76.68 57.32 24.55 74.13 67.19 35.11 25.42 62.96 21.47 49.43 11.44 SpectralGPT 80.47 57.99 35.44 89.07 33.42 37.85 46.95 58.86 26.75 51.87 10.11 S.-S12-MoCo 81.58 51.76 34.49 89.26 53.02 35.44 48.58 57.64 25.38 53.02 10.06 S.-S12-DINO 81.72 49.37 36.18 88.61 51.15 34.81 48.66 56.47 25.62 52.51 10.89 S.-S12-MAE 81.91 49.90 32.03 87.79 51.92 34.08 45.80 57.13 24.69 51.69 12.39 S.-S12-Data2Vec 81.91 44.36 34.32 88.15 48.82 35.90 54.03 58.23 24.23 52.22 10.72 UNet Baseline 84.51 54.79 31.60 91.42 60.47 39.46 47.57 62.09 46.34 57.58 4.89 ViT Baseline 81.58 48.19 38.53 87.66 59.32 36.83 44.08 52.57 38.37 54.13 10.28 TerraMindv1-B-single 84.00 65.01 40.80 90.32 52.66 59.71 27.71 TerraMindv1-B 82.42 69.52 40.51 90.62 59.72 37.87 55.80 60.61 28.12 58.35 3.94 TerraMindv1-L 82.93 75.57 43.13 90.78 63.38 37.89 55.04 59.98 27.47 59.57 3.44
Table 11: Performance evaluation of TerraMind across seven benchmark datasets using the PANGAEA evaluation protocol. Higher mIoU values ($\uparrow$) and lower rank values ($\downarrow$) are reported. The best model per column is highlighted in bold and the second best is underscored. We indicate unimodal datasets with , demonstrating that TerraMind outperforms other benchmarks in both unimodal and multimodal settings.
Since the first approaches to foundation models for Earth observations, experts in the field discuss on the usability of such models compared to task-specific models that are trained for each application individually. Recent benchmark results suggested that task-specific models, like U-Nets, often outperform finetuned GFMs [49]. We therefore additionally investigate how TerraMind compares with task-specific U-Nets and ViT models following the PANGAEA evaluation protocol. As adviced by the authors of PANGAEA, we again report results on nine of the eleven datasets as we could not reproduce the performance on the remaining two datasets. The task-specific models are trained from scratch for each individual task, while all GFMs including TerraMind are finetuned with a frozen encoder and an UperNet head. For reasons of consistency, we report TerraMind results based on full finetuning instead of TiM tuning. Yet, for several tasks, we expect slightly higher results when applying TiM tuning for TerraMind. Overall, our results demonstrate that TerraMindv1-B outperforms task-specific UNet and ViT models across the PANGAEA benchmark in both unimodal and multimodal settings by 1pp avg. mIoU and 4pp avg. mIoU respectively. In multimodal settings, the improvement peaks to 4.5pp improvement of TerraMindv1-B over task-specific U-Nets. To the best of our knowledge, this is the first time a GFM model outperforms task-specific models on a global benchmark.
In addition, we observe that for most datasets, TerraMindv1-B outperforms TerraMindv1-B-single. This demonstrates the benefit from scaling in the data and feature dimension–i.e., leveraging dual-scale feature representations on a pixel level and a token level.
12.3 Comparing generations and pseudo-labels
We evaluate the model generations for modalities where we used pseudo-labels as input data. For example, in initial experiments with TerraMind.v0.1, we leverage Google’s DynamicWorld model to pseudo-label LULC maps which we use as input to the model. In the following experiment in Table 12, we test the performance of the DynamicWorld model against the generations of TerraMind. Overall, we observe that while finetuned TerraMind.v0.1-B outperforms DynamicWorld, the generation of TerraMind does not surpass the inference results of DynamicWorld.
Approach Input IoUWater
TerraMindv1-B-single S-2 L1C 69.87
Dynamic World pseudo-labeling S-2 L1C 71.98
TerraMindv1-B-single finetuning S-2 L1C 76.32
Table 12: Results on the Sen1Floods11 test set comparing flood maps derived from TerraMind’s out-of-the-box LULC generations to those derived from LULC pseudo-labeling with Dynamic World. The results are inferior to those obtained by fine-tuning a specialized model for this downstream task, which is expected.
12.4 TiM tuning for crop mapping
We further investigate the relevance of TiM tuning for crop type mapping in order to understand the relevance of generating artificial data for more finegrained segmentation tasks. That means, we generate artificial LULC data which includes agricultural crop as a single class and investigate whether this additional information helps to segment nine different types of crops in satellite images. We experiment with the South Africa Crop Type Mapping dataset (https://source.coop/esa/fusion-competition) and present the results in Table 13. Overall, we observe that TiM tuning improves the performance by around 1pp. That means that even though the generated artificial data does not include further information on the location and shape of certain crops, the information on where to expect crop land in general helps to guide the model to an improved performance.
Input mIoU
TerraMindv1-B S-2 41.87
TerraMindv1-B TiM S-2 + gen. LULC 42.74
Table 13: Thinking-in-modalities (TiM) tuning compared with standard full fine-tuning approaches on the SA Crop dataset.